OpenStudy (anonymous):

The profit P (in hundreds of dollars) that a company makes depends on the amount x (in hundreds of dollars) the company spends on advertising according to the model shown below. P = 260 + 70x – 0.5x^2. What expenditure for advertising will yield a maximum profit?

OpenStudy (anonymous):

differentiate P with respect of x \[\frac{{\rm d}P}{{\rm d}x} = 70 - x\] set the derivative equal to zero \[ 70 - x = 0 \] \[x = 70 \] that in hundreads of dollars, so result is $7000.